OpenStudy (anonymous):

find the number of sides of a regular polygon with each interior angle equal to 160 degrees?

OpenStudy (anonymous):

exterior angle = 180 - 160 = 20 No of sides = 360 / 20 = 18

OpenStudy (anonymous):

by interior angle u mean the angle made at the centre ?

OpenStudy (anonymous):

inside the polygon

OpenStudy (amistre64):

interior angles are: 360 - exterior

OpenStudy (amistre64):

160/2 = 80 80+80+central angle = 180 central angle = 20 360/#sides = 20 360/20 = #sides = 18 ...yep